The lengths of the three sides of a triangle add to 35. One side is eight more than two times as long as the smallest. The third side is eight less than seven times as long as the smallest. What is the measure of each side? (Enter your answers as a comma-separated list.)
Answer by JulietG(1812)

A + b + c = 35
Let C be the third side
Let B be the smallest side
Let A be "one side"
A = 2B + 8 [One side is eight more than two times as long as the smallest.]
C = 7B - 8 [The third side is eight less than seven times as long as the smallest.]
Now take the known values of A & C and substitute them into the top equation.
A + B + C = 35
(2B+8) + B + (7B-8) = 35
10B = 35
B = 3.5
If B is 3.5, then A is 15, and C is 16.5
3.5 + 15 + 16.5 = 35
The sides are 15,3.5,16.5